# Aberdeen Proving Ground, USA

### v2

@@ 4.2.2 Geometry of Pareto optimality
$d_{avg}(f,g)\approx\frac{2}{3}\frac{d^{3}_{max}-d^{3}_{min}}{d^{2}_{max}-d^{2}_{min}}.$ (31)
@@
The average deviation from conjugacy over all possible Pareto optimal solutions is obtained by converting to polar coordinates via $r_{1}=\gamma\cos\theta$ and $r_{2}=\gamma\sin\theta$, and then computing
@@ 4.3 Analytical solutions in $\Phi$-space
While an optimal permutation matrix $P$ is one solution to Eq. 37 in general it is not the solution to the conjugacy residual problem given by Eq. 28; see Remark 4. Since any solution $C_{r_{2}}=\Gamma P$ such that $\sum_{i}|\lambda_{f,i}-\lambda_{g,i}|^{2}$ is minimized will minimize Eq. 37, one also requires a solution for $\Gamma$. Thus a more general solution than a pure permutation matrix is required for computing deviation from conjugacy. The solution for $\Gamma$ proposed here is obtained by finding the best solution (in a least squares sense) that minimizes Eq. 35. Substituting $C=\Gamma P$ into Eq. 35 gives
@@ 4.4 Solutions in $\Psi$-space
While the primary focus here is computing deviation from conjugacy pseudometrics, one may be interested in analyzing systems in $\Psi$-space. Therefore, for the sake of completeness transformations in $\Psi$-space are recovered by substituting Eqs. 36 and 40 into Eq. 14. Since solutions for $T$ are based on left-eigenvectors which are not unique, an optimal magnitude for the eigenvectors must be found. Recognizing that $W_{g}$ can be arbitrarily scaled by a diagonal matrix $\Omega$, Eq. 14 becomes
@@ 5.1 Benchmarking example
Consider the systems $f$ and $g$ given by Eqs. 45 and 46 respectively.
@@
Data is generated for $\lambda=-0.1+10i$ and $\mu=0.001+i$ and a uniform grid over the parameters $\alpha$ and $\beta$ by varying them from 0.1 to 2 in increments of 0.01. Discrete time data is generated for $N_{\Delta t}=1000$ time steps using a discretization of $\Delta t=0.01$. At conjugacy, the non-orthogonal transformation in $\Psi$-space between the two systems is
@@
Additional insight can be gained through visualization of the trajectories. For instance consider the two cases illustrated in Fig. 5. The operator residual $r_{2}(C_{r_{2}})$ for both cases in Figs. 5(a) and 5(b) is 0.20. For reference, the least squares solution $T_{LSQ}=\Psi_{g}\Psi_{f}^{\dagger}$ which is the best possible solution in $GL(N_{\Psi})$ from a trajectory geometry perspective,is provided. In the case shown in Figs. 5(a), the optimal operator based solution $C_{r_{2}}$and optimal trajectory geometry solution $C_{r_{2}}$ lead to nearly identical results. In this case, it would not matter whether conclusions about dissimilarity are made by using $r_{1}$ or $r_{2}$. In contrast, Fig. 5(b) illustrates a case in which the optimal solutions for $r_{1}$ or $r_{2}$ are different. The trajectory residuals are $r_{1}(C_{r_{1}})=1.03$ and $r_{1}(C_{r_{2}})=1.28$ for this case. So while accounting for $r_{2}$ alone may lead one to conclude that the cases $\alpha=1,\beta=1.2$ and $\alpha=1.85,\beta=0.94$ are equally dissimilar to the case $\alpha,\beta=1$, accounting for both $r_{1}$ and $r_{2}$ instead indicates that the system corresponding to $\alpha=1.85,\beta=0.94$ is more dissimilar to the $\alpha,\beta=1$ system compared to $\alpha=1,\beta=1.2$. This is confirmed in the deviation from conjugacy pseudmoetrics which are $d_{min}\approx d_{avg}\approx d_{max}=0.58$ for the $\alpha=1,\beta=1.2$ case and $d_{min}=1.05$, $d_{avg}=1.18$, and $d_{max}=1.31$ for the $\alpha=1.85,\beta=0.94$ case.
@@ 5.2 Engineering example
As an engineering example, morphological computation pfeifer_bongard of a single degree of freedom mechanical actuator driven by a controller to achieve stable hopping is considered. Morphological computation is a measure of how much the body can offload the required information processing of the controller. Systems with high morphological computation will require less information to be processed by a control system in order to achieve a task compared to a system with low morphological computation. In particular, actuation and control models representative of biological muscle and DC motor actuators are compared MCt ; haeufle_models . The quantification of state-dependent morphological computation from MCt is used because the dynamics of morphological computation were shown to provide more insight than time averaged quantities. Since morphological computation is a state-dependent quantity it can be treated as a Koopman observable.
@@
NLM/LM 0.91 1.12 1.31 0.91 0.06 1.31 0.01
NLM/DC 0.29 0.73 1.03 0.29 0.05 1.03 0.01
LM/DC 0.94 1.22 1.46 0.94 0.06 1.46 0.02
The deviation from conjugacy pseudometrics and associated conjugacy residuals are provided in Table 1. The pseudometrics are calculated using normalized residuals from Eqs. 19 and 20, with the nonlinear muscle actuator serving as the reference system. The deviations from conjugacy quantify the intuitive notion that all of the dynamical systems are quite different. However, the finding that the DC motor attempt to approximate biological muscle is closer than the linearized muscle approximation for all pseudometrics is interesting and perhaps not intuitive. Furthermore, it is interesting to note that the linearized muscle and DC motor systems are very different from each other with $d_{avg}=1.22$ even though their time-averaged morphological computation values are relatively close; 4.76 bits for the linearized muscle and 4.52 bits for the DC motor compared to 7.30 bits for the nonlinear muscle. As a result, one may conclude based on mean values of morphological computation that the linearized muscle and the DC motor actuator are relatively similar approximations of nonlinear muscle. However a comparison based on the dynamically consistent pseudometrics capture differences due to dynamics of morphological computation that may be lost when comparing systems based on time-averaged morphological computation. Such comparisons reveal that the DC motor is relatively dissimilar to the linearized muscle.
@@
Figures 10 and 11 provide visual insight to the numerical values in Table 1. Table 1 shows the tradeoffs between minimizing $r_{1}$ versus $r_{2}$ for these systems. The DC motor system is moderatley dissimilar to the nonlinear muscle system when considering trajectory geometry as indicated by $r_{1}(C_{r_{1}})=0.29$. Optimizing for $r_{2}$ leads to relatively modest improvements from $r_{2}(C_{r_{1}})=0.05$ to $r_{2}(C_{r_{2}})=0.01$. However, this comes at the cost of increasing $r_{1}$ to 1.03. This is ilustrated in Fig. 10 where the transformation based on optimizing for $r_{1}$ shows that the transformed nonlinear muscle system resembles the $\Psi_{\mathrm{primary}}$ geometry of the DC motor system. Interestingly, the transformed solution in Fig. 10(b) corresponding to minimizing $r_{2}$ resembles the shape of the DC motor $\Psi_{\mathrm{primary}}$ trajectories, but does not match the scale. For comparison, Fig. 11 shows that the nonlinear muscle system does not resemble the linearized muscle system when transformed.

### v4

@@ 4.2.2 Geometry of Pareto optimality
$\begin{split}&d_{avg}(f,g)\approx\sqrt{(r_{1}(C_{r_{1}})+\Delta_{r_{1}}/2)^{2}+(r_{2}(C_{r_{2}})+\Delta_{r_{2}}/2)^{2}}\\ &+[\Delta_{r_{2}}^{3}\sinh^{-1}(\Delta_{r_{1}}/\Delta_{r_{2}})+\Delta_{r_{1}}^{3}\sinh^{-1}(\Delta_{r_{2}}/\Delta_{r_{1}})\\ &+2\Delta_{r_{1}}\Delta_{r_{2}}\sqrt{\Delta_{r_{1}}^{2}+\Delta_{r_{2}}^{2}}]/(12\Delta_{r_{1}}\Delta_{r_{2}}),\end{split}$ (31)
where
$\Delta_{r_{1}}=r_{1}(C_{r_{2}})-r_{1}(C_{r_{1}})$ (32)
and
$\Delta_{r_{2}}=r_{2}(C_{r_{1}})-r_{2}(C_{r_{2}}).$ (33)
@@
The average deviation from conjugacy over all points in the rectangle depicted in Fig. 1 is given by the integral
@@
The integral can be split into the sum of two parts $d_{avg}=d_{avg,1}+d_{avg,2}$: the first part is given by the distance from the origin to the center of the rectangle depicted in Fig. 1, and second is the average distance from all points in the rectangle to the rectangle’s center. The first part, denoted by $d_{avg,1}$ follows directly from the geometry of Fig. 1 and is given by
$d_{avg,1}=\sqrt{(r_{1}(C_{r_{1}})+\Delta_{r_{1}}/2)^{2}+(r_{2}(C_{r_{2}})+\Delta_{r_{2}}/2)^{2}},$ (37)
where $\Delta_{r_{1}}$ and $\Delta_{r_{2}}$ are the lengths of the rectangle’s sides.
The second part $d_{avg,2}$ follows from rectangle_dist and is given by
$\begin{split}&d_{avg,2}=[\Delta_{r_{2}}^{3}\sinh^{-1}(\Delta_{r_{1}}/\Delta_{r_{2}})+\Delta_{r_{1}}^{3}\sinh^{-1}(\Delta_{r_{2}}/\Delta_{r_{1}})\\ &+2\Delta_{r_{1}}\Delta_{r_{2}}\sqrt{\Delta_{r_{1}}^{2}+\Delta_{r_{2}}^{2}}]/(12\Delta_{r_{1}}\Delta_{r_{2}})\end{split}$ (38)
@@ 4.3 Analytical solutions in $\Phi$-space
While an optimal permutation matrix $P$ is one solution to Eq. 41, in general it is not the solution to the conjugacy residual problem given by Eq. 28; see Remark 4. Since any solution $C_{r_{2}}=\Gamma P$ such that $\sum_{i}|\lambda_{f,i}-\lambda_{g,i}|^{2}$ is minimized will minimize Eq. 41, one also requires a solution for $\Gamma$. Thus a more general solution than a pure permutation matrix is required for computing deviation from conjugacy. The solution for $\Gamma$ proposed here is obtained by finding the best solution (in a least squares sense) that minimizes Eq. 39. Substituting $C_{r_{2}}=\Gamma P$ into Eq. 39 gives
@@ 4.4 Solutions in $\Psi$-space
While the primary focus here is computing deviation from conjugacy pseudometrics, one may be interested in analyzing systems in $\Psi$-space. Therefore, for the sake of completeness transformations in $\Psi$-space are recovered by substituting Eqs. 40 and 44 into Eq. 14. Since solutions for $T$ are based on left-eigenvectors which are not unique, an optimal magnitude for the eigenvectors must be found. Recognizing that $W_{g}$ can be arbitrarily scaled by a diagonal matrix $\Omega$, Eq. 14 becomes
@@ 5.1 Benchmarking example
Consider the systems $f$ and $g$ given by Eqs. 49 and 50 respectively.
@@
Data is generated for $\lambda=-0.1+10i$ and $\mu=-0.001+i$ and a uniform grid over the parameters $\alpha$ and $\beta$ by varying them from 0.1 to 2 in increments of 0.01. Discrete time data is generated for $N_{\Delta t}=1000$ time steps using a discretization of $\Delta t=0.01$. At conjugacy, the non-orthogonal transformation in $\Psi$-space between the two systems is
@@
Additional insight can be gained through visualization of the trajectories. For instance consider the two cases illustrated in Fig. 5. The operator residual $r_{2}(C_{r_{2}})$ for both cases in Figs. 5(a) and 5(b) is 0.20. For reference, the least squares solution $T_{LSQ}=\Psi_{g}\Psi_{f}^{\dagger}$ which is the best possible solution in $GL(N_{\Psi})$ from a trajectory geometry perspective,is provided. In the case shown in Figs. 5(a), the optimal operator based solution $C_{r_{2}}$and optimal trajectory geometry solution $C_{r_{1}}$ lead to nearly identical results. In this case, it would not matter whether conclusions about dissimilarity are made by using $r_{1}$ or $r_{2}$. In contrast, Fig. 5(b) illustrates a case in which the optimal solutions for $r_{1}$ and $r_{2}$ are different. The trajectory residuals are $r_{1}(C_{r_{1}})=1.03$ and $r_{1}(C_{r_{2}})=1.28$ for this case whereas the residuals for the case depicted in Fig. 5(a) are $r_{1}(C_{r_{1}})=r_{1}(C_{r_{2}})=0.55$. So while accounting for $r_{2}$ alone may lead one to conclude that the cases $\alpha=1,\beta=1.2$ and $\alpha=1.85,\beta=0.94$ are equally dissimilar to the case $\alpha,\beta=1$, accounting for both $r_{1}$ and $r_{2}$ instead indicates that the system corresponding to $\alpha=1.85,\beta=0.94$ is more dissimilar to the $\alpha,\beta=1$ system compared to $\alpha=1,\beta=1.2$. This is confirmed in the deviation from conjugacy pseudmoetrics which are $d_{min}=d_{avg}=d_{max}=0.58$ for the $\alpha=1,\beta=1.2$ case and $d_{min}=1.05$, $d_{avg}=1.24$, and $d_{max}=1.31$ for the $\alpha=1.85,\beta=0.94$ case.
@@ 5.2 Engineering example
As an engineering example, morphological computation pfeifer_bongard of a single degree of freedom mechanical actuator driven by a controller to achieve stable hopping is considered. Morphological computation is a measure of how much the body can offload the required information processing of the controller. Systems with high morphological computation will require less information to be processed by a control system in order to achieve a task compared to a system with low morphological computation. In particular, actuation and control models representative of biological muscle and DC motor actuators are compared MCt ; haeufle_models . The quantification of state-dependent morphological computation from MCt is used because the dynamics were shown to provide more insight than time averaged quantities. Since morphological computation is a nonlinear state-dependent quantity it can be treated as a Koopman observable.
@@
NLM/LM 0.91 1.13 1.20 0.91 0.06 1.20 0.01
NLM/DC 0.29 0.72 0.86 0.29 0.05 0.86 0.01
LM/DC 0.94 1.23 1.32 0.94 0.06 1.32 0.02
The deviation from conjugacy pseudometrics and associated conjugacy residuals are provided in Table 1. The pseudometrics are calculated using normalized residuals from Eqs. 19 and 20, with the nonlinear muscle actuator serving as the reference system. The deviations from conjugacy quantify the intuitive notion that all of the dynamical systems are quite different. However, the finding that the DC motor attempt to approximate biological muscle is closer than the linearized muscle approximation for all pseudometrics is interesting and perhaps not intuitive. Furthermore, it is interesting to note that the linearized muscle and DC motor systems are very different from each other with $d_{avg}=1.23$ even though their time-averaged morphological computation values are relatively close; 4.76 bits for the linearized muscle and 4.52 bits for the DC motor compared to 7.30 bits for the nonlinear muscle. As a result, one may conclude based on mean values of morphological computation that the linearized muscle and the DC motor actuator are relatively similar approximations of nonlinear muscle. However a comparison based on the dynamically consistent pseudometrics capture differences due to dynamics of morphological computation that may be lost when comparing systems based on time-averaged morphological computation.
@@
Table 1 shows the tradeoffs between minimizing $r_{1}$ versus $r_{2}$ for these systems. Figures 10 and 11 provide visual insight to theses numerical values. The DC motor system is moderatley dissimilar to the nonlinear muscle system when considering trajectory geometry as indicated by $r_{1}(C_{r_{1}})=0.29$. Optimizing for $r_{2}$ leads to relatively modest improvements from $r_{2}(C_{r_{1}})=0.05$ to $r_{2}(C_{r_{2}})=0.01$. However, this comes at the cost of increasing $r_{1}$ to 0.86. This is illustrated in Fig. 10 where the transformation based on optimizing for $r_{1}$ shows that the transformed nonlinear muscle system resembles the $\Psi_{\mathrm{primary}}$ geometry of the DC motor system. Interestingly, the transformed solution in Fig. 10(b) corresponding to minimizing $r_{2}$ resembles the shape of the DC motor $\Psi_{\mathrm{primary}}$ trajectories, but does not match the scale. For comparison, Fig. 11 shows that the nonlinear muscle system does not resemble the linearized muscle system when transformed, which tracks with the higher values reported in Table 1.